OpenStudy (lukecrayonz):

y=2(x-10)^2+1 How do I convert to standard form?

Directrix (directrix):

Convert y=2(x-10)^2+1 to the y = ax^2 + bx + c form. First, crank out and simplify (x-10)^2 Multiply what you get by 2 and then add 1 to the sum. @Lukecrayonz

OpenStudy (lukecrayonz):

How did you see this when it was closed o.O

Directrix (directrix):

Closed threads are visible by clicking on the Closed Questions tab to the right of the Open Questions tab. This has always been the case.

OpenStudy (lukecrayonz):

Literally have never noticed that. Thank you though :-) 2x^2−40x+201

Directrix (directrix):

Don't forget the y. y = 2x^2−40x+201

OpenStudy (lukecrayonz):

Who created openstudy?:O

OpenStudy (lukecrayonz):

I've been on here for about two years and you've been here the whole time, so I'm assuming you played a big role in coding?

OpenStudy (lukecrayonz):

Just found out, Ashwin Ram! Didn't know Openstudy was funded by the National Science Foundation either.. impressive!